Evaluate the 5/9+9/8?

Answer by Boreal(15235)

Common denominator is 72:
5/9=40/72, multiplying top and bottom by 8
9/8=81/72, multiplying top and bottom by 9
Now they can be added over the same denominator of 72
The answer is 121/72 or 1 49/72